OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

slope = (5-(-4))/(2-(-1)) = 9/3 =3 equation: y-5 = 3(x-2) it doesn't matter which poing you use in the final answer. the reason why i chose (2,5) is purely subjective.

OpenStudy (anonymous):

slope =3 y-5=3(x-2) y-5=3x-6 3x-y-6+5= 3x-y-1= this is the equation